OpenStudy (anonymous):

Write an Equation of the line that passes through the given point and has the given slope: "(3,1) Slope 2"

OpenStudy (anonymous):

I need the answer w/ work.

OpenStudy (anonymous):

ok, i'll help do you remember the slope intercept form of y=mx+b?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

so if we were to have an equation of a line, then that means we will use the format of y=mx+b in the question we are given: x, y and m, so we need to figure out what 'b' is inorder to write the equation. to find 'b' replace all the variables in the y=mx+b equation with the numbers we know, then solve for 'b'

OpenStudy (anonymous):

So y=3x+2?

OpenStudy (anonymous):

1x*

OpenStudy (anonymous):

close, that is how the answer should look when we are done, with only the x and y showing and numbers for the rest, but what I meant to say was solve for 'b' if x=3 y=1 m=2 in y=mx+b

OpenStudy (anonymous):

Ah. Okay. Carry on

OpenStudy (anonymous):

so, then we'd have 1=2(3)+b do you see it? solve for 'b' now ^_^

OpenStudy (anonymous):

@Happymail123 still need help or you got it? i'm just hanging around to help..

OpenStudy (anonymous):

What would the answer be in the end?

OpenStudy (anonymous):

y=2x+b, but you have to plug in b. b is from this equation i put up before, 1=2(3)+b

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

yep ^_^ have a good day